Circular No. 7650

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

V1178 SCORPII N. N. Samus and E. Kazarovets, Institute of Astronomy, Moscow, in-

form us that the designation V1178 Sco has been given to the novalike variable announced on IAUC 7647.

G. Masi, Ceccano, Italy, reports that his CCD observations obtained show V1178 Sco at unfiltered mag 11.1 on June 21.934 UT, based on USNO-A2.0 red magnitudes for the stars in the field. Astrometry yields the following position for the variable: α = 17^{h}57^{m}06^{s}.73, δ = −32^{o}23′05′′.6 (equinox 2000.0; eight USNO-A2.0 reference stars). Masi adds that there is a USNO-A2.0 star having position end figures 06^{s}.71, 05′′.5 with red mag 16.5 and blue mag 17.2; another star very close to this has position end figures 06^{s}.63, 05′′.6 with red mag 15.2 and blue mag 15.6;

Additional magnitude estimates: Apr. 18.749 UT, [11.2 (K. Kanatsu, Matsue, Shimane, Japan, T-Max 400 film; communicated by T. Kato, Kyoto University); 26.735, [10.6 (Kanatsu); May 13.673, 10.3 (Kanatsu); 19.672, [10.6 (Kanatsu); June 21.980, 10.81 (W. Liller, Viña del Mar, Chile, broadband V); 22.108, 10.87 (Liller); 22.173, 10.89 (Liller); 22.245, 10.97 (Liller); 22.249, 10.96 (Liller).

SUNGRAZING COMETS Further to IAUC 7646, D. Hammer reports his measurements for six

additional Kreutz sungrazing comets on images obtained with the LASCO C2 coronagraph onboard SOHO; comet C/2001 M3 was also visible in C3 images. Comet C/2001 M6 showed a short tail on a few images. The comets were found on SOHO website images by X. Leprette (C/1997 W3), M. Oates (C/2001 M2, M4), S. Hoenig (C/2001 M3), X. Zhou (C/2001 M5), and T. Scarmato (C/2001 M6). The reduced astrometry and orbital elements by B. G. Marsden appear on MPEC 2001-M33.

Comet  1997 UT  α_{2000}  δ_{2000}
C/1997 W3 Nov. 24.775  16^{h}01.^{m}4  −22^{o}16′
Comet  2001 UT  α_{2000}  δ_{2000}
C/2001 M2 June 16.594  5^{h}36.^{m}2  +21^{o}34′
C/2001 M3  16.684  5 37.2  +21 18
C/2001 M4  19.013  5 46.3  +21 39
C/2001 M5  20.271  5 51.1  +21 42
C/2001 M6  21.996  5 57.8  +21 43
2001 June 23  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green